sqrt x+7 + sqrt x+16=9
thank you

Answer by stanbon(75887)

sqrt(x+7) + sqrt(x+16) = 9
Square both sides to get:
(x+7)+(x+16)+2sqrt(x^2+23x+112)=81
2x+23+ sqrt(------------)=81
sqrt(---------) = -2x+58
------------
Square both sides again to get:
x^2+23x+112 = 4x^2-232x+3364
3x^2-255x+3252=0
x^2-85x+1084=0
x=[85+-sqrt(7225-4*1084)]/2
x=[85+-sqrt(2889)]/2
x=[85+53.75]/2=69.37
or
x=[85-53.75]/2=15.625
-------
Cheers,
Stan H.